Lillian is about to invest $20,000, part at 6% and the rest at 8%.
What is the most that she can invest at 6% and still be gauranteed at least $1500 in intrest per year?

Answer by Fermat(136)

Let Lillian invest $P @ 6%
Let Lillian also invest $(20,000 - P) @ 8%
Total revenue is:
P*6/100 + (20,000 - P)*8/100 >= 1500
6P + 160,000 - 8P >= 150,000
-2P >= -10,000
2P <= 10,000
P <= 5,000
==========
Thus Lillian must invest no more than $5,000 @ 6% (with the remainder @ 8%) inorder to ensure that her total revenue will exceed $1500.